OpenStudy (anonymous):

The total snowfall in Burke, Virginia, this January was 7 1/4 in. In February, 20% more snow fell than in January. How much snow fell in February in Burke, Virginia? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

@Agl202

OpenStudy (pawanyadav):

Complete question first...

OpenStudy (anonymous):

what?

OpenStudy (pawanyadav):

Read the question ....

Vocaloid (vocaloid):

the question looks fine to me... first convert 7 1/4 to a decimal

OpenStudy (anonymous):

7.25?

Vocaloid (vocaloid):

good now, to find a 20% increase, we multiply that by 1.2 7.25*1.2 = ?

OpenStudy (anonymous):

8.7? @Vocaloid

OpenStudy (pawanyadav):

20% increase mean Initial +20% of initial rainfall So. 7.25 + 20% of 7.25 = 7.25 + 0.2*7.25

Vocaloid (vocaloid):

good, now convert that back into a mixed number

OpenStudy (anonymous):

wow i forgot how to do that ;-;

OpenStudy (pawanyadav):

Very good....

OpenStudy (pawanyadav):

Then leave it in decimal

OpenStudy (pawanyadav):

This is also correct

OpenStudy (pawanyadav):

Do you want to learn ...conversion in mixed number

Vocaloid (vocaloid):

whoops, sorry for poofing to convert 8.7 to a mixed number, let's just look at the 0.7 part first 0.7 is the same as 7/10 8.7 = 8 7/10